A custodian is going to mix a 4% ammonia solution and a 12% ammonia solution to get 1 gallon (128 fluid ounces) of a 9% ammonia solution. How many fluid ounces of the 4% solution and the 12% solution should be used?
Answer by richwmiller(17219)

We need a total of 128 oz. at 9%
We want to know how many oz. at 4% to mix with 12%
a+b=128,
0.04*a+0.12*b=0.09*128
a=128-b
0.04*(128-b)+0.12*b=11.52
5.12-0.04b+0.12*b=11.52
0.08*b=6.4
b=80
a=128-b
a=48 oz. at 4%
b=80 oz. at 12%
check
0.04*48+0.12*80=0.09*128
1.92+9.6=11.52
11.52=11.52
ok
codemx1